# Can Language Models Recognize Convincing Arguments?

### v1

@@ 5 Results
Considering the PoliIssues dataset [IS], we summarize the accuracy of the different LLMs and baseline methods in correlating beliefs and demographic characteristics with stances on Table 2 (rows #8–#15). Here the accuracy range of different LLMs is much more narrow, ranging from 41.39% (Mistral) to 42.82% (GPT-4). Most important however, is that the performance of LLMs is similar to that of crowdworkers (39.32%; MTurk).
@@
Llama False False 41.30 (37.92, 44.38)
@@
We study whether the results obtained were sensitive to prompt used by re-running the analysis from RQ2 on the PoliIssues dataset. For each model, we rerun the analysis considering the “big issues” in user profiles and/or asking for models to reason before answering. [^1] The prompt constraint was changed to Evaluate step-by-step the data given in the proposition before coming to an answer. Provide your reasoning for selecting an answer and then give your answer in the form of ‘Pro,’ ‘Con,’ or ‘Tie’ without using other words or punctuation. Provide your response in the following format: ‘Reasoning: your reasoning goes here. Answer: your answer goes here.’ Results are shown in Table 3. Overall, we find that the results are not sensitive to the experimented changes.
@@
6 Discussion
@@
Our findings indicate that large language models demonstrate human-level performance in (1) judging argument quality, (2) predicting users’ stances on specific topics given users’ demographics and basic beliefs, and (3) detecting arguments that would be persuasive to individuals with specific demographics or beliefs. However, the overall human performance is not high in each of the three tasks [around 60% for (1), and around 40% for (2) and (3)], which could be due to the inherent difficulty of the tasks, as well as variance and randomness in the data. This does not necessarily imply that LLMs do not pose any additional risk of tailor-made misinformation in the future. It is plausible that with access to more personal information about an individual, such as personality traits, LLMs could perform better at detecting persuasive arguments (Hirsh et al., 2012). Nevertheless, it is important to consider that the more fine-grained the target, the harder and more costly it becomes to reach the targeted population, and the cost-benefit analysis is not straightforward (Tappin et al., 2023).
@@
One hypothesis that could explain the relatively low accuracy for both LLMs and human performance is that these demographic questions and big-issue stances may not be highly relevant for the task, as suggested by Hu & Collier (2024). However, this is contradicted by the fact that a supervised XGBoost model trained with these factors yields much better results. Interestingly, stacking various LLM predictions yields performance closer to XGBoost. This indicates that while an individual LLM may not excel at detecting persuasive arguments for an individual, combining the predictions of several LLMs could achieve much more competitive performance. Consequently, LLMs can potentially generate highly effective tailored misinformation and propaganda, particularly in a multi-agent setting Schoenegger et al. (2024).
@@
One limitation of our work is that the ability to detect personalized persuasive content, while closely related, is not equivalent to the ability to generate such content. However, we argue that since generation is easier than detection for LLMs (West et al., 2023), our work contributes a valuable and efficient framework for continuously assessing the persuasive capabilities of LLMs. Another limitation is that our dataset, from debate.org, may not be representative of the general population. The demographics of individuals opting to participate in online debates are likely skewed compared to the U.S. population and even more so globally. Additionally, we could not test the language models on non-English data due to data access limitations. However, recent research has shown that language models’ performance is considerably lower for non-English languages, especially low-resource ones (Ahuja et al., 2023). Consequently, it is plausible that the risk of misuse for microtargeting in non-English settings is currently lower. Nevertheless, as language models continue to improve, it is crucial to expand this line of research to a wider range of languages and demographics to ensure a comprehensive understanding of the risks associated with personalized persuasion.
Ethics.
@@
In this study, we employ demographic and belief-related questions drawn from datasets that are publicly accessible and have been anonymized before release. It is crucial to emphasize the importance of responsible development and deployment of LLMs and the need for ongoing research into mitigating their potential risks Bommasani et al. (2021). Our work can inform the development of safeguards and countermeasures against the misuse of LLMs for personalized misinformation and propaganda. Finally, in this work, we used crowdsourcing to rate the persuasiveness of debates. We paid crowd workers (based in the US) \$12.00/hour median rate; higher than the US federal minimum wage.
Reproducibility.
We release our code for modeling and analysis. [^2] Code: https://anonymous.4open.science/r/debate-gpt-x-88F3/ [^3] Data: https://osf.io/8jh62/?view_only=c8c28914ca664330b02cfd476e097abc We include details about the crowdsourcing setup in Appendix B.
@@ Appendix A Prompts
Figure 4: Prompt structure used in RQ2.
Figure 5: Prompt structure used in RQ3.
In all tasks conducted for this study, the LLMs were prompted to respond to questions using only one of the options: ”Pro,” ”Con,” or ”Tie,” without any additional words or punctuation. Nevertheless, many of the models failed to adhere to this instruction, necessitating post-processing to extract the actual answer. Generally, incorrect responses involved the answer being accompanied by additional spaces or punctuation or presented in a complete sentence format, such as: “Based on the given demographics, the person is most likely to agree with the ‘Con’ side in the debate.” [^4] There were occasions when the LLMs failed to produce any answer, resulting in responses akin to: “I cannot determine the person’s position in the debate without additional information.” To extract the answer in these cases, we used a simple Regex expression, finding the first occurrence of the words “Pro,” “Con,” or “Tie.”
@@
Table 4: Some models had difficulty following instructions and giving the answer in the correct form of either ”Pro,” ”Con,” or ”Tie.” In this table, we see what percentage of the answers were given in the correct form and what percentages contained an answer after processing the result for the PoliProp dataset.
Appendix B Crowdsourcing
@@
  • Disregarding your own point of view on the debate, please determine which debater you believe had more convincing arguments. The individual arguing for the proposition (Pro) or against it (Con)? If both were similarly convincing, indicate that it was a ”Tie.”

### v2

@@ 5 Results
Considering the PoliIssues dataset [IS], we summarize the accuracy of the different LLMs and baseline methods in correlating beliefs and demographic characteristics with stances on Table 2 (rows #8–#15). Here, the accuracy range of different LLMs is much more narrow, ranging from 41.39% (Mistral) to 42.82% (GPT-4). Most important, however, is that the performance of LLMs is similar to that of crowdworkers (39.32%; MTurk).
@@
Llama 2 False False 41.30 (37.92, 44.38)
@@
We study whether the results obtained were sensitive to the prompt used by re-running the analysis from RQ2 on the PoliIssues dataset. For each model, we rerun the analysis considering the “big issues” in user profiles and/or asking for models to reason before answering. [^1] The prompt constraint was changed to Evaluate step-by-step the data given in the proposition before coming to an answer. Provide your reasoning for selecting an answer and then give your answer in the form of ‘Pro,’ ‘Con,’ or ‘Tie’ without using other words or punctuation. Provide your response in the following format: ‘Reasoning: your reasoning goes here. Answer: your answer goes here.’ Results are shown in Table 3. Overall, we find that the results are not sensitive to the experimented changes.
@@
6 Discussion and Conclusion
Here we studied LLM’s persuasive capabilities by considering its ability to identify convincing arguments in general and for people with specific arguments. We argue that if LLMs can detect content that is highly persuasive to specific demographics, they may be used to detect and amplify tailor-made misinformation and propaganda. Our findings indicate that LLMs demonstrate human-level performance in (1) judging argument quality, (2) predicting users’ stances on specific topics given users’ demographics and basic beliefs, and (3) detecting arguments that would be persuasive to individuals with specific demographics or beliefs.
@@ 6 Discussion and Conclusion
However, the overall human performance is not high in each of the three tasks [around 60% for (1), and around 40% for (2) and (3)], which could be due to the inherent difficulty of the tasks, as well as variance and randomness in the data. This does not necessarily imply that LLMs do not pose any additional risk of tailor-made misinformation in the future. It is plausible that with access to more personal information about an individual, such as personality traits, LLMs could perform better at detecting persuasive arguments (Hirsh et al., 2012). Nevertheless, it is important to consider that the more fine-grained the target, the harder and more costly it becomes to reach the targeted population, and the cost-benefit analysis is not straightforward (Tappin et al., 2023).
@@
One hypothesis that could explain the relatively low accuracy for both LLMs and human performance is that these demographic questions and big-issue stances may not be highly relevant for the task, as suggested by Hu and Collier (2024). However, this is contradicted by the fact that a supervised XGBoost model trained with these factors yields much better results. Interestingly, stacking various LLM predictions yields performance closer to XGBoost. This indicates that while an individual LLM may not excel at detecting persuasive arguments for an individual, combining the predictions of several LLMs could achieve much more competitive performance (perhaps because each LLM’s biases differ). Consequently, LLMs can potentially detect highly effective tailored misinformation and propaganda, particularly in a multi-agent setting Schoenegger et al. (2024).
Limitations
@@
Our dataset, from debate.org, may not be representative of the general population. The demographics of individuals opting to participate in online debates are likely skewed compared to the U.S. population and even more so globally. Additionally, we could not test the language models on non-English data due to data access limitations. However, recent research has shown that language models’ performance is considerably lower for non-English languages, especially low-resource ones (Ahuja et al., 2023). Consequently, it is plausible that the risk of misuse for microtargeting in non-English settings is currently lower. Nevertheless, as language models continue to improve, it is crucial to expand this line of research to a wider range of languages and demographics to ensure a comprehensive understanding of the risks associated with personalized persuasion. It is also essential to conduct empirical studies to understand whether LLMs are, in fact, being used for persuasion in online settings (e.g., in social media platforms). Another limitation of the work at hand is that the LLMs studied might have seen content from debate.org in their training data. To address this concern, we queried 100 debate excerpts from the dataset using GPT4 and couldn’t obtain complete samples. Yet, this is not sufficient to rule out this possibility.
Ethical Considerations
@@
In this study, we employ demographic and belief-related questions drawn from datasets that are publicly accessible and have been anonymized before release. It is crucial to emphasize the importance of responsible development and deployment of LLMs and the need for ongoing research into mitigating their potential risks Bommasani et al. (2021). Our work can inform the development of safeguards and countermeasures against the misuse of LLMs for personalized misinformation and propaganda.
In this study, we employed crowdsourcing to evaluate the persuasiveness of debates. We paid crowd workers, all based in the U.S., at a rate of \$12.00 per hour, higher than the federal minimum wage in the United States.
@@ Appendix A Prompts
In all tasks conducted for this study, the LLMs were prompted to respond to questions using only one of the options: "Pro," "Con," or "Tie," without any additional words or punctuation. Nevertheless, many of the models failed to adhere to this instruction, necessitating post-processing to extract the actual answer. Generally, incorrect responses involved the answer being accompanied by additional spaces or punctuation or presented in a complete sentence format, such as: “Based on the given demographics, the person is most likely to agree with the ‘Con’ side in the debate.” [^2] There were occasions when the LLMs failed to produce any answer, resulting in responses akin to: “I cannot determine the person’s position in the debate without additional information.” To extract the answer in these cases, we used a simple Regex expression, finding the first occurrence of the words “Pro,” “Con,” or “Tie.”
@@
Table 4: Some models had difficulty following instructions and giving the answer in the correct form of either "Pro," "Con," or "Tie." In this table, we see what percentage of the answers were given in the correct form and what percentages contained an answer after processing the result for the PoliProp dataset.
Appendix B Demographics and Big Issues
The dataset by Durmus and Cardie (2018) contained the following demographic information about participants: birthday, education, ethnicity, gender, income, party, political ideology, religious ideology.
It also contained participants’ opinions on so-called “big issues.” They were: abortion, affirmative action, animal rights, Barack Obama, border fence, capitalism, civil unions, death penalty, drug legalization, electoral college, environmental protection, estate tax, European Union, euthanasia, federal reserve, flat tax, free trade, gay marriage, global warming exists, globalization, gold standard, gun rights, homeschooling, Internet censorship, Iran-Iraq war, labor union, legalized prostitution, Medicaid and medicare, medical marijuana, military intervention, minimum wage, national health care, national retail sales tax, occupy movement, progressive tax, racial profiling, redistribution, smoking ban, social programs, social security, socialism, stimulus spending, term limits, torture, United Nations, war in Afghanistan, war on terror, and welfare.
Appendix C Crowdsourcing
@@
  • Disregarding your own point of view on the debate, please determine which debater you believe had more convincing arguments. The individual arguing for the proposition (Pro) or against it (Con)? If both were similarly convincing, indicate that it was a "Tie."